# Spectrum of random centrosymmetric matrices; CLT and Circular law

### v2

@@ 7.3. Double chain Expectation
For this case, we need to impose more constraints that will reduce the degree of freedom and leads to zero asymptotic contribution. Here, $x_{i_{1}i_{2}}=\bar{x}_{j_{3}j_{4}}$ if and only if $(i_{1},i_{2})=(j_{3},j_{4})$ or the dual case $(i_{2},i_{3})=(n+1-j_{3},n+1-j_{4})$.
If $j_{3}=i_{1}$, this implies $j_{4}=i_{2}$. Furthermore, since we are pairing $x_{i_{2}i_{3}}$ and $\bar{x}_{j_{5}j_{1}}$, we shall have $j_{5}=i_{2}$ and $i_{3}=j_{1}$, or the dual case. Now, let us look at the consecutive random variables of those already paired up. We observe that to maintain the continuity of the chain, we must impose restrictions that will significantly reduce the number of free indices and lead to zero asymptotic contribution.
In conclusion, the total number of combinations that provide a nonzero asymptotic contribution is $2\times 5n^{5}$. We obtain that contribution from the cases as demonstrated in the Figure 4. Consequently,
@@
Let us expand $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{k})]$ as follows;
$\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{k})]$ $=\frac{1}{n^{k}}\mathbb{E}\left[\sum_{{i_{1},i_{2},\dots i_{k}\atop j_{1},j_{2},\dots j_{k}=1}}^{n}(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\cdots\bar{x}_{j_{k}j_{1}})\right]$
$=\frac{1}{n^{k}}\sum_{{i_{1},i_{2},\dots i_{k}\atop j_{1},j_{2},\dots j_{k}=1}}^{n}\mathbb{E}\left[(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\cdots\bar{x}_{j_{k}j_{1}})\right].$
@@
The main contribution will come from two possible cases, such as (a) the chain $i_{1}\rightarrow i_{2}\rightarrow\dots\rightarrow i_{k}\to i_{1}$ overlaps completely with the chain $j_{1}\rightarrow j_{2}\rightarrow\dots\rightarrow j_{k}\to j_{1}$, or (b) it overlaps with the dual version of the second chain, which is $n+1-j_{1}\rightarrow n+1-j_{2}\rightarrow\dots\rightarrow n+1-j_{k}\rightarrow n+1-j_{1}.$ In either cases, we have $k$ many possible ways to start the second chain. Once the chains overlap, it reduces to the product of terms of the form $\mathbb{E}[|x_{ij}|^{2}]$ $(=1)$. Such complete overlapping will contribute $2kn^{k}.$
@@
Now, we argue that if the chains do not overlap completely, then the contribution is negligible. Let us assume that only $p$ random variables from the the first chain overlap with $p$ random variables in the second chain. In that case, the overlapping part contributes $2pn^{p}$. Additionally, the extra $k-p$ random variables from the first and the second chain may contribute at most $n^{(k-p-1)/3}(k-p)^{3(k-p)}$ each. This is a consequence of the result stated in the Lemma 7.4 with a chain of length $k-p$. Note that in the Lemma 7.4, we had the factor $n^{-k/6}$, where $k$ was the length of the chain, because the trace sum was already divided by $n^{k/2}$. Here, length of the chain is $k-p$, but the factor $n^{-(k-p)/2}$ is not taken into account yet. Consequently, we would get a contribution of $n^{(k-p)/3}$. However, to keep the continuity of the full chain (e.g. $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k}i_{1}}$), one of the index among $k-p$ variables is already determined by the remaining $p$ variables which are matched with $p$ variables of the another chain (e.g. $x_{j_{1}j_{2}}x_{j_{2}j_{3}}\cdots x_{j_{k}j_{1}}$). As a result, we have the factor $n^{(k-p-1)/3}.$ Hence in such a case, the combined contribution turns out to be at most
(7.1) $2pn^{p}n^{2(k-p-1)/3}(k-p)^{6(k-p)}$ $=n^{k}\times 2pn^{-(k-p+2)/3}(k-p)^{6(k-p)}.$
@@
Additionally, the starting points of the $p$-chains can be set to one of indices among $i_{1},i_{2},\ldots,i_{k}$ (for the first chain) or $j_{1},j_{2},\ldots,j_{k}$ (for the second chain). Therefore, by multiplying a factor of $k^{2}$ and summing up over all possible values of $p$, we see that if the full chains do not overlap, then we have the following contribution
(7.2) $n^{k}\sum_{p=1}^{k-3}2pk^{2}n^{-(k-p+2)/3}(k-p)^{6(k-p)}.$
@@
Note that $p$ can not be $k-1$ or $k-2$. Because in that case, the remaining chain would be of length one or two, and from the Condition 3.3, we have $\mathbb{E}[x_{ij}]=0=\mathbb{E}[x_{ij}^{2}].$ The above sum is $o(n^{k})$ if $k<n^{\gamma}$ for $\gamma<1/20.$ Thus, we have
@@
$=\frac{1}{n^{k}}\sum_{\begin{subarray}{c}i_{1},i_{2},\dots i_{k},\\ j_{1},j_{2},\dots j_{k}=1\end{subarray}}^{n}\mathbb{E}\left[(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\cdots\bar{x}_{j_{k}j_{1}})\right]$
@@
Without loss of generality, let us assume that $k<l$. We adopt the methodology from Lemma 7.5. Let a sub chain of length $p$ from the first chain $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\cdots x_{i_{k}i_{1}}$ gets matched to a sub chain of length $p$ from the second chain $x_{j_{1}j_{2}}x_{j_{2}j_{3}}\cdots x_{j_{l}j_{1}}.$ In that case, following the equations (7.1) and (7.2), we have
@@
While calculating $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$, the significant contribution comes only when $k=l$, see Lemma 7.5. From Lemma 7.4 and Lemma 7.6, we conclude that if $k\neq l$, although intra-chain merging and partial-chain merging are possible strategies, asymptotically they provide zero contributions. In particular, $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]=O(|l-k|^{3|l-k|}n^{-|l-k|/6})$ for $k,l\leq n^{1/20}$.
@@
$=\sum_{k,l=1}^{d}a_{k}\bar{a}_{l}\operatorname{Cov}\left(\operatorname{Tr}(M^{k}),\operatorname{Tr}(M^{l})\right)$
$=\sum_{k,l=1}^{d}a_{k}\bar{a}_{l}\left(\mathbb{E}\left[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})\right]-\mathbb{E}\left[\operatorname{Tr}(M^{k})\right]\mathbb{E}\left[\operatorname{Tr}(\bar{M}^{l})\right]\right)$
(7.3) $=\sum_{\begin{subarray}{c}k\neq l\\ k,l=1\end{subarray}}^{d}0+\sum_{k=1}^{d}2k|a_{k}|^{2}+o(1)$
$=\sum_{k=1}^{d}2k|a_{k}|^{2}+o(1).$
@@ 7.4. Covariance kernel of $\operatorname{Tr}\hat{\mathcal{R}}_{z}(M)$
We see that on the event $\Omega_{n}$, the last term is bounded by $\frac{n}{\tau}(\rho/|z|)^{\lfloor\log^{2}n\rfloor}=o(1).$ Hence, using Remark 7.7, we have
@@
As a result, $\sum_{k\neq l=1}^{\lfloor\log^{2}n\rfloor}z^{(-k-1)}\bar{z}^{(-l-1)}\mathbb{E}[\operatorname{Tr}M^{k}\operatorname{Tr}\bar{M}^{l}]=o(1).$ Hence, by using Lemma 7.5, we obtain
@@
=$=\left\{\mathbb{E}[\operatorname{Tr}\hat{\mathcal{R}}_{z}(M)\operatorname{Tr}\hat{\mathcal{R}}_{\bar{\eta}}(\bar{M})]-\frac{n^{2}}{z\bar{\eta}}\right\}+o(1)$
@@
[…]
@@
In our case, we have a matrix with i.i.d. entries whose entries satisfy the Poincaré inequality. Therefore, by taking the variance profile $\omega\equiv 1$, and using the above lemma we can have the following.
@@
where $K>0$ is a universal constant, which does not depend on the size of the matrix and the bandwidth.

### v3

@@ 7.3. Double chain Expectation
For this case, we need to impose more constraints that will reduce the degree of freedom and leads to zero asymptotic contribution. Here, $x_{i_{1}i_{2}}=\bar{x}_{j_{3}j_{4}}$ if and only if $(i_{1},i_{2})=(j_{3},j_{4})$ or the dual case $(i_{2},i_{3})=((n+1)-j_{3},(n+1)-j_{4})$.
If $j_{3}=i_{1}$, this implies $j_{4}=i_{2}$. Furthermore, since we are pairing $x_{i_{2}i_{3}}$ and $\bar{x}_{j_{5}j_{1}}$, we shall have $j_{5}=i_{2}$ and $i_{3}=j_{1}$, or the dual case. Now, look at the consecutive random variables of those already paired up. We observe that to maintain the continuity of the chain, we must impose restrictions that will significantly reduce the number of free variables indices and lead to zero asymptotic contribution.
Therefore, the total number of combinations that provide a nonzero asymptotic contribution is $2\times 5n^{5}$. We obtain that contribution from the cases as demonstrated in the Figure 4. Consequently,
@@
The lemma estimates the limiting expectation of cross chain merging. Consider
$\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{k})]$ $=\frac{1}{n^{k}}\mathbb{E}\left[\sum_{{i_{1},i_{2},\dots i_{k}\atop j_{1},j_{2},\dots j_{k}=1}}^{n}(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\dots\bar{x}_{j_{k}j_{1}})\right]$
$=\frac{1}{n^{k}}\sum_{{i_{1},i_{2},\dots i_{k}\atop j_{1},j_{2},\dots j_{k}=1}}^{n}\mathbb{E}\left[(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\dots\bar{x}_{j_{k}j_{1}})\right].$
@@
The main contribution will come from two possible cases, such as (a) the chain $i_{1}\rightarrow i_{2}\rightarrow\dots\rightarrow i_{k}\to i_{1}$ overlaps completely with the chain $j_{1}\rightarrow j_{2}\rightarrow\dots\rightarrow j_{k}\to j_{1}$, or (b) it overlaps with the dual version of the second chain, which is $(n+1-j_{1})\rightarrow(n+1-j_{2})\rightarrow\dots\rightarrow(n+1-j_{k})\rightarrow(n+1-j_{1}).$ In either cases, we have $k$ many possible ways to start the second chain. Once the chains overlap, it reduces to the product of terms of the form $\mathbb{E}[|x_{ij}|^{2}]$ $(=1)$. Such complete overlapping will contribute $2kn^{k}.$
@@
Now we argue that if the chains do not overlap completely, then the contribution is negligible. Let us assume that only $p$ random variables from the the first chain overlap with $p$ random variables in the second chain. In that case, the overlapping part contributes $2pn^{p}$. Additionally, the extra kpk-p random variables from the first and the second chain may contribute at most $n^{(k-p-1)/3}(k-p)^{3(k-p)}$ each. This a consequence of the result stated in the Lemma 7.4. Note that in the Lemma 7.4, we had the factor $n^{-k/6}$, where $k$ was the length of the chain, and the trace sum was already divided by $n^{k/2}$. Here length of the chain is kpk-p, but it is not divided by $n^{(k-p)/2}$ yet. Consequently, we are supposed to get $n^{(k-p)/3}$ instead of $n^{-(k-p)/6}.$ However, to keep the continuity of the full chain (e.g. $x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}}$), one of the index among kpk-p variables is already determined by the remaining $p$ variables which are matched with $p$ variables of the another chain (e.g. $x_{j_{1}j_{2}}x_{j_{2}j_{3}}\dots x_{j_{k}j_{1}}$). As a result, we have the factor $n^{(k-p-1)/3}.$ Hence in such a case, the combined contribution turns out to be at most
(7.1) $2pn^{p}n^{2(k-p-1)/3}(k-p)^{6(k-p)}$ $=n^{k}\times 2pn^{-(k-p-2)/3}(k-p)^{6(k-p)}.$
@@
Additionally, the starting points of the $p$-chains can be set to one of indices among $i_{1},i_{2},\ldots,i_{k}$ (for the first chain) or $j_{1},j_{2},\ldots,j_{k}$ (for the second chain). Therefore multiplying by a factor of $k^{2}$ and summing up over all possible values of $p$, we see that if the full chains do not overlap, then we have the following contribution
(7.2) $n^{k}\sum_{p=1}^{k-3}2pk^{2}n^{-(k-p-2)/3}(k-p)^{6(k-p)}.$
@@
Note that $p$ can not be k1k-1 or k2k-2. Because in that case, the remaining chain would be of length one or two, and from the Condition 3.3, we have $\mathbb{E}[x_{ij}]=0=\mathbb{E}[x_{ij}^{2}].$ The above sum is $o(n^{k})$ if $k<n^{\gamma}$ for $\gamma<1/20.$ Therefore, we have
@@
$=\frac{1}{n^{k}}\sum_{\begin{subarray}{c}i_{1},i_{2},\dots i_{k},\\ j_{1},j_{2},\dots j_{k}=1\end{subarray}}^{n}\mathbb{E}\left[(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\dots\bar{x}_{j_{k}j_{1}})\right]$
@@
[…]
@@
Proof.
Here the chain lengths are not equal. Without loss of generality, let us assume that $k<l$. We can divide the second chain into two parts - one with the length of the first chain and one with the remaining elements, as follows; $(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}})$ $(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\dots\bar{x}_{j_{k}j_{k+1}})$ $(\bar{x}_{j_{k+1}j_{k+2}}\bar{x}_{j_{k+2}j_{k+3}}\dots\bar{x}_{j_{l}j_{1}})$. To get the maximum contribution, the first and second chains will have cross chain merging, and the third chain will have intra chain merging.
From the cross chain merging of the first and second chains, we have $2kn^{k}$ combinations; see the Lemma 7.5. Now, we have one option for the third chain, which is intra chain merging. We have at most $n^{(l-k-1)/3}(l-k)^{3(l-k)}$ combinations from intra chain merging. Note that the random variable $j_{k+1}$ has already been fixed in the previous case where we merged the first and second chains. Additionally, we chose the second and third chains from a single chain, and this can be done in $l$ many ways. Therefore, we have at most $ln^{(l-k-1)/3}(l-k)^{3(l-k)}$ combinations. Thus, the total number of combinations is of order $O(lk(l-k)^{3(l-k)}n^{(l+2k-1)/3})$. However, when evaluating $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$, the expression $\mathbb{E}\left[(x_{i_{1}i_{2}}x_{i_{2}i_{3}}\dots x_{i_{k}i_{1}})(\bar{x}_{j_{1}j_{2}}\bar{x}_{j_{2}j_{3}}\dots\bar{x}_{j_{l}j_{1}})\right]$ is multiplied by $(1/n)^{(k+l)/2}$. Therefore, from this case we have a contribution of $O((l-k)^{3(l-k)}n^{-(l-k)/6})$. ∎
@@
While calculating $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]$, the significant contribution comes only when $k=l$, see Lemma 7.5. From Lemma 7.4 and Lemma 7.6, we conclude that if $k\neq l$, although intra-chain merging and partial-chain merging are possible strategies, asymptotically they provide zero contributions. In particular, $\mathbb{E}[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})]=O((l-k)^{3(l-k)}n^{-(l-k)/6})$.
@@
$=\sum_{k,l=1}^{d}a_{k}a_{l}\operatorname{Cov}\left(\operatorname{Tr}(M^{k}),\operatorname{Tr}(M^{l})\right)$
$=\sum_{k,l=1}^{d}a_{k}a_{l}\left(\mathbb{E}\left[\operatorname{Tr}(M^{k})\operatorname{Tr}(\bar{M}^{l})\right]-\mathbb{E}\left[\operatorname{Tr}(M^{k})\right]\mathbb{E}\left[\operatorname{Tr}(\bar{M}^{l})\right]\right)$
(7.3) $=\sum_{\begin{subarray}{c}k\neq l\\ k,l=1\end{subarray}}^{d}0+\sum_{\begin{subarray}{c}k=l\\ k,l=1\end{subarray}}^{d}2ka_{k}^{2}+o(1)$
$=\sum_{k=1}^{d}2ka_{k}^{2}+o(1).$
@@ 7.4. Covariance kernel of $\operatorname{Tr}\hat{\mathcal{R}}_{z}(M)$
We see that on the event $\Omega_{n}$, the last term is bounded by $\frac{n}{\tau}(\rho/|z|)^{\lfloor\log^{2}n\rfloor}=o(1).$ Now using Remark 7.7, we have
@@
As a result, $\sum_{k\neq l=1}^{\lfloor\log^{2}n\rfloor}z^{(-k-1)}\bar{z}^{(-l-1)}\mathbb{E}[\operatorname{Tr}M^{k}\operatorname{Tr}\bar{M}^{l}]=o(1).$ Therefore using Lemma 7.5, we obtain
@@
$=\left\{\mathbb{E}[\operatorname{Tr}\hat{\mathcal{R}}_{z}(M)\operatorname{Tr}\hat{\mathcal{R}}_{\bar{\eta}}(\bar{M})]-\frac{n^{2}}{z\bar{\eta}}\right\}+o(1)$
@@
[17, Lemma A.3] Let $M$ be a random band matrix with bandwidth $b_{n}$ and variance profile $\omega$. Assume that the entries are i.i.d. having mean zero and variance one, and satisfy the Poincaré inequality with constant $\alpha.$ Then there exists $\rho\geq 1$ such that
@@
In our case, we have a matrix with i.i.d. entries whose entries satisfy the Poincaré inequality. Therefore, using the above lemma we can have the following.
@@
where $K>0$ is a universal constant.
@@
Acknowledgment
Indrajit Jana’s research is partially supported by INSPIRE Fellowship
DST/INSPIRE/04/2019/000015, Dept. of Science and Technology, Govt. of India.